### v1

# Meta-Statistical Learning: Supervised Learning of Statistical Inference

Abstract
This work demonstrates that the tools and principles driving the success of large language models (LLMs) can be repurposed to tackle distribution-level tasks, where the goal is to predict properties of the data-generating distribution rather than labels for individual datapoints. These tasks encompass statistical inference problems such as parameter estimation, hypothesis testing, or mutual information estimation. Framing these tasks within traditional machine learning pipelines is challenging, as supervision is typically tied to individual datapoint. We propose meta-statistical learning, a framework inspired by multi-instance learning that reformulates statistical inference tasks as supervised learning problems. In this approach, entire datasets are treated as single inputs to neural networks, which predict distribution-level parameters. Transformer-based architectures, without positional encoding, provide a natural fit due to their permutation-invariance properties. By training on large-scale synthetic datasets, meta-statistical models can leverage the scalability and optimization infrastructure of Transformer-based LLMs. We demonstrate the framework’s versatility with applications in hypothesis testing and mutual information estimation, showing strong performance, particularly for small datasets where traditional neural methods struggle.
Machine Learning, ICML
1 Introduction
Statistical inference is the backbone of many scientific inquiries, providing a rigorous framework for quantifying evidence, testing hypotheses, and estimating uncertainty (Walker and Lev, 1953; Casella and Berger, 2024). Across scientific disciplines, statistical inference is an accepted mechanism through which scientists relate noisy observations to theoretical models; it underpins the design of experiments, the validation of theories, and the interpretation of empirical results (Barlow, 1993; Altman, 1990; James, 2006; Dienes, 2008; Salganik, 2019).
However, the practice of statistical inference is notoriously difficult. Real-world data is noisy often deviating from idealized assumptions (Gurland and Tripathi, 1971; Hoekstra et al., 2012; Knief and Forstmeier, 2021; Czyż et al., 2023). In particular, inference in low-sample regimes presents a persistent challenge, yet it is crucial across many applied sciences. In such settings, estimators must balance universality with bias and variance, whereas more robust estimators require strong assumptions on the underlying distribution (Casella and Berger, 2024). Some statistical quantities simply lack universally unbiased estimators – like the standard deviation – necessitating context-dependent correction strategies (Gurland and Tripathi, 1971; Bengio and Grandvalet, 2003). In general, designing statistical estimators requires making choices regarding the bias-variance and robustness-universality trade-offs, requiring manual effort to craft estimators to specific goals (Silvey, 2013).
Machine learning, itself a form of statistical inference, can provide a flexible approach to these challenges. Instead of manually designing statistical estimators, we propose to learn them from data, leveraging amortized learning strategies to train models that generalize across diverse data distributions and adapt their estimation strategy contextually to new inputs. We call this approach meta-statistical learning, wherein entire datasets are treated as input objects and statistical inference tasks are directly framed as supervised learning problems (see Figure 1).
Figure 1: Illustration of the Meta-Statistical learning setup: a meta-distribution $P_{\Gamma}$ dictates the sampling of meta-datapoints, couples of datasets $X$ and the label $y$ a property of data-generating distribution $P_{X}$. The meta-statistical model learns to predict $y$ from entire datasets effectively converting statistical inference in a supervised learning problem.
Meta-statistical learning shifts the unit of analysis from individual data points to entire datasets. Unlike traditional supervised learning, where the goal is to predict a label $y$ for an individual sample $x$ drawn from a joint distribution $P_{X,Y}$, meta-statistical models learn to map datasets to their target statistical properties from large amounts of synthetic datasets. This formulation aligns naturally with modern deep learning tools, which can easily handle various input modalities such as images (Voulodimos et al., 2018), graphs (Ma and Tang, 2021), time-series (Gamboa, 2017; Lim and Zohren, 2021; Torres et al., 2021), language (Vaswani et al., 2017; Devlin et al., 2019; Radford et al., 2019), and tabular data (Gulati and Roysdon, 2023; Hollmann et al., 2025). The approach remains firmly within supervised learning where datasets are just another input modality processed by neural networks. However, this perspective enables us to flexibly tackle statistical inference tasks by supervised learning.
The success of large language models (LLMs) (Vaswani et al., 2017; Radford et al., 2019; Achiam et al., 2023) serves as both an inspiration and a blueprint for the meta-statistical research direction. First, attention-based architectures inherently satisfy the permutation invariance property necessary for processing datasets as unordered collections (Lee et al., 2019; Zhang et al., 2022), making them default candidate architectures. Second, the substantial infrastructure developed for LLMs—including algorithmic frameworks, hardware optimizations, and software ecosystems—can be readily repurposed for meta-statistical modeling. Finally, large-scale training datasets can be synthetically generated for most statistical inference tasks, enabling robust generalization across data distributions and strong performance in low-sample settings where classical estimators often struggle.
Contributions. In this work: (i) We introduce the meta-statistical framework and establish its connections to related paradigms such as meta-learning and amortized inference (Section 2). (ii) We evaluate various meta-statistical architectures, demonstrating that dataset encoders based on Set Transformer (Lee et al., 2019) variants achieve strong performance and generalization while avoiding the quadratic computational cost of standard attention mechanisms (Section 4). (iii) We showcase the versatility of the framework through three statistical inference problems: estimating standard deviation, conducting normality tests, and estimating mutual information. Our results highlight the robustness and generalization capabilities of meta-statistical models across diverse data distributions and, especially in low-sample-size settings (Section 5). (iv) We release for practitioners meta-statistical normality test and MI estimator, strong in low-sample regimes.
We believe the meta-statistical framework offers a promising path to leverage the principles that made LLMs successful, re-purposing them to tackle challenging statistical inference problems.
2 Meta-Statistical Learning
2.1 Background: Supervised Learning
Supervised learning aims to find a function $f:\mathcal{X}\to\mathcal{Y}$ that maps input data to output labels based on a finite dataset of observations $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}$ and $i\in\{1,\dots,n\}$. Here, $\mathcal{X}$ denotes the input space (e.g., $\mathbb{R}^{d}$ for $d$-dimensional data), and $\mathcal{Y}$ denotes the output space, which is continuous for regression or discrete for classification. The data points are assumed to be i.i.d. samples from an unknown joint distribution $P_{X,Y}$ over $\mathcal{X}\times\mathcal{Y}$.
The function $f$ is modeled by a parameterized family $\{f_{\theta}:\theta\in\Theta\}$, where $\theta$ represents the parameters (e.g., weights in a neural network). The quality of $f_{\theta}$ is evaluated using a loss function $L:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$, which measures the discrepancy between predicted and true outputs.
Generalization. The goal is to minimize the expected risk: $R(\theta)=\mathbb{E}_{(x,y)\sim P_{X,Y}}[L(f_{\theta}(x),y)]$, but since $P_{X,Y}$ is unknown, the empirical risk is minimized as a proxy. Generalization is achievable because machine learning algorithms perform induction, based on assumptions about the underlying structure of the data and the expectation of how new data relate to observed ones. Typically, we expect the model to generalize in distribution, where new instances are sampled from $P_{X,Y}$. However, we often also care about generalization out of distribution, where new instances are sampled from a different, but related distribution.
2.2 Meta-Statistical Learning
Instead of learning a mapping from individual data points to their labels, meta-statistical learning maps entire datasets to their labels. Meta-statistical learning remains within standard supervised learning with the dataset being just another modality representable by a neural network.
Setup and notation. Meta-statistical learning aims to find a function $\varphi:\Gamma\to\mathcal{Y}$ that maps input datasets to labels based on a finite meta-dataset $\mathcal{S}=\{(\mathcal{D}_{i},y_{i})\}_{i=1}^{M}$, where $\mathcal{D}_{i}\in\Gamma$ is itself a dataset $\mathcal{D}_{i}=\{(x_{i,j})\}_{j=1}^{n_{i}}$. As in standard supervised learning, $\mathcal{Y}$ denotes the output space, which is continuous for regression or discrete for classification. The meta-datapoints are assumed to be sampled i.i.d. from an unknown joint meta-distribution $P_{\Gamma,Y}$, a distribution over datasets (their data-generating distribution) and their target labels. The function $\varphi$ is modeled by a parameterized family $\{\varphi_{\theta}:\theta\in\Theta\}$ that can process entire datasets as input (e.g., a recurrent neural network, convolutional neural network, or Transformer). The quality of $\varphi_{\theta}$ is still evaluated using a loss function $L_{\Gamma}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$. The learning objective remains to minimize the expected risk, but taken over the meta-distribution:
$R(\theta)=\mathbb{E}_{(\mathcal{D},y)\sim P_{\Gamma,Y}}\left[L_{\Gamma}\left(\varphi_{\theta}(\mathcal{D}),y\right)\right].$
2.3 Structure of the Meta-Generalization Problem
To provide additional structure to the generative process that produces a meta-datapoint $(\mathcal{D}_{i},y_{i})\sim P_{\Gamma,Y}$, we decompose it into two steps: (i) sample a distribution $P_{X}$, and (ii) sample a dataset $\mathcal{D}_{i}\sim P_{X}$. This process is illustrated in Figure 1. The label can either be a property of the dataset itself, $y_{i}=A(\mathcal{D}_{i})$, or a property of the distribution, $y_{i}=g(P_{X})$. When the label is a property $A$ of the dataset, we refer to it as a descriptive label, such as the column-wise average. When the label is a property $g$ of the distribution, we refer to it as an inferential label, such as determining whether the dataset was sampled from a normal distribution or estimating the mutual information between two variables.
Several generalization questions arise from this setup:
(i) Within-distribution generalization: The function $\varphi_{\theta}$ should generalize across different datasets resampled from the same distribution $P_{X}$. In the inferential case, where the label depends only on $P_{X}$, $\varphi_{\theta}$ should produce the same prediction for all datasets of fixed size sampled from $P_{X}$. The predictions of $\varphi_{\theta}$ should not systematically overestimate or underestimate the label. This is measured by the variance and the bias of $\varphi_{\theta}$ as a statistical estimator of $y=g(P_{X})$.
(ii) Length generalization: The function $\varphi_{\theta}$ should generalize to datasets of varying lengths. Statistical inference is harder for smaller datasets, so we expect performance to improve with larger datasets. This is measured by the consistency of $\varphi_{\theta}$ as a statistical estimator of $y=g(P_{X})$.
(iii) In-meta-distribution generalization: Similar to standard supervised learning, $\varphi_{\theta}$ should generalize to new meta-datapoints sampled from the same meta-distribution $P_{\Gamma,Y}$. For example, if $\varphi_{\theta}$ is trained to predict the standard deviation of datasets sampled from exponential distributions, it should generalize to exponential distributions with unseen rate parameters.
(iv) Out-of-meta-distribution generalization: Analogous to out-of-distribution generalization, $\varphi_{\theta}$ could be expected to generalize to distributions and datasets sampled from a different meta-distribution than $P_{\Gamma,Y}$. For instance, if $\varphi_{\theta}$ is trained on datasets from Normal, Uniform, and Exponential distributions, it can be tested on datasets sampled from Log-normal, Cauchy, or Weibull distributions.
2.4 Related Work
The idea of processing multiple data points simultaneously originates from multi-instance learning, where models receive sets of instances and assign labels at the group level (Maron and Lozano-Pérez, 1997; Dietterich et al., 1997; Ilse et al., 2018). Once datasets can be meaningfully represented by neural networks, amortized learning techniques allowing models to generalize quickly to new datasets naturally emerge (Ganguly et al., 2023; Lopez-Paz et al., 2015; Kim et al., 2024).
A notable example is the neural statistician framework (Edwards and Storkey, 2017), which employs variational autoencoders (VAEs) to learn dataset representations in an unsupervised manner. Similarly, Hewitt et al. (2018) applied VAEs to infer generative models from few data points. The concept of learning dataset-level representations has also been explored through meta-features (Jomaa et al., 2021; Kotlar et al., 2021; Hartmann et al., 2023), where models extract high-level statistics tailored for specific tasks. For instance, Kotlar et al. (2021) learned meta-features for anomaly detection, while Wu et al. (2022) trained models to predict dataset-level statistics such as the number of distinct values. Recently, Hollmann et al. (2025) employed transformers trained on synthetic datasets for missing value imputation, which we recognize as an instance of meta-statistical learning in low-sample-size settings.
Approaches of a meta-statistical nature have also been successfully applied in causal discovery (Lopez-Paz et al., 2015; Löwe et al., 2022; Lorch et al., 2022; Wu et al., 2024). These methods generate synthetic data with known causal structures and train neural networks to infer causal properties from a set of observations (Ke et al., ). For example, Kim et al. (2024) proposed an attention-based model trained on simulated datasets to identify causal parents of target variables. Meta-statistical learning is a type of amortized learning focused on estimating statistical parameters; it builds upon and generalizes these previous works.
Relationship to Meta-Learning. Meta-learning, or learning to learn, is a paradigm focused on generalizing across tasks drawn from different distributions (Schmidhuber et al., 1996; Hospedales et al., 2021; Huisman et al., 2021). Meta-learning seeks to acquire transferable meta-knowledge, enabling rapid adaptation to new tasks (Schmidhuber, 1987; Thrun, 1998; Schmidhuber, 1993; Vanschoren, 2019). A broad range of approaches exist (Vinyals et al., 2016; Santoro et al., 2016; Finn et al., 2017; Snell et al., 2017), some emphasizing dataset-level processing to extract useful representations (Mishra et al., 2017; Ravi and Larochelle, 2017; Munkhdalai and Yu, 2017; Shyam et al., 2017). This is particularly relevant in few-shot learning (Finn et al., 2017; Snell et al., 2017; Wang et al., 2023; Wu et al., 2020; Rivolli et al., 2022). Notably, neural processes represent a class of meta-learners that use a meta-distribution over functions, adapting their prior to new datasets using observed input-output pairs (Garnelo et al., 2018b, a; Kim et al., 2019). Meta-statistical learning shares conceptual similarities with meta-learning, as both focus on generalization across distributions. However, while the target of meta-learning remains instance-level predictions, meta-statistical learning emphasizes distributional properties. These paradigms are complementary: insights from dataset-level analysis can directly improve generalization in meta-learning (Jomaa et al., 2021; Kotlar et al., 2021).
3 Experimental Setup
Our experiments demonstrate the versatility of meta-statistical learning by achieving strong performance across diverse tasks with minimal task-specific effort. Here, we describe the template used to run experiments with various descriptive and inferential tasks.
Meta-Dataset Generation. We construct meta-datasets by sampling datasets $\mathcal{D}$ and labels $y$ from a predefined meta-distribution $P_{\Gamma,Y}$. The generation process involves two stages: first, a distribution family (e.g., Normal, Uniform) is randomly selected, and its parameters are sampled from predefined priors to yield a data-generating distribution $P_{X}$. A dataset $\mathcal{D}$ of size $n$ is then sampled from $P_{X}$, with $n$ also drawn from a prior. Thus, $P_{\Gamma}$ defines the set of base distributions, parameter priors, and dataset size priors.
In- vs. Out-of-Meta-Distribution. With in-meta-distribution (IMD) settings, both training and testing datasets are sampled from $P_{\Gamma}$. For out-of-meta-distribution (OoMD) testing, we modify $P_{\Gamma}$ by changing the set of base distributions (e.g., replacing Normal with Cauchy). This tests the robustness of meta-statistical estimators to unseen distributions.
3.1 Meta-Statistical Models
Models should predict dataset-level properties $y$ from datasets $\mathcal{D}$ of varying sizes $n$. The architecture we consider consists of a dataset encoder $\phi$ and a prediction head $\rho$, defined as $\varphi(\mathcal{D};\rho,\phi)=\rho\circ\phi(\mathcal{D})$, where $\phi$ transforms $\mathcal{D}$ into a fixed-dimensional representation, and $\rho$ is a Multi-Layer Perceptron that predicts the target. The model is trained with MSE loss for regression tasks and cross-entropy loss for classification tasks.
LSTM Encoder. The Long Short-Term Memory (LSTM) network (Hochreiter and Schmidhuber, 1997) processes datasets sequentially. For a dataset $\mathcal{D}\in\mathbb{R}^{n\times k}$, the dataset representation is the average of all hidden states of the last layer: $\phi(\mathcal{D})=\frac{1}{n}\sum_{t=1}^{n}h_{t}$, where $h_{t}$ is the hidden state at timestep $t$. LSTMs lack the inductive bias of permutation invariance, making them a baseline model.
Vanilla Transformer Encoder. The Vanilla Transformer (VT) (Vaswani et al., 2017) uses multi-head self-attention without positional encodings to ensure permutation invariance. The dataset representation is the output of a special token, analogous to the CLS token in BERT (Devlin et al., 2019): $\phi(\mathcal{D})=z_{\text{CLS}}.$
Set Transformer. The Set Transformer (Lee et al., 2019) is designed for set-structured data and ensures permutation invariance. Furthermore, it reduces the quadratic cost of attention by performing attention on a fixed set of $m$ inducing points, where $m$ is a hyperparameter. The inducing points are learned as a projection of the full sequence at each layer. The enhanced Set Transformer 2 (ST2) (Zhang et al., 2022) incorporates SetNorm, a normalization technique that improves over LayerNorm (Ba, 2016) by preserving permutation invariance while improving the convergence properties of the Set Transformer.
4 Experiments on Descriptive Tasks
In descriptive tasks, the label $y$ of a dataset $\mathcal{D}$ is the output of an algorithm $A$ applied to $\mathcal{D}$, i.e., $y=A(\mathcal{D})$. Simple tasks like median or correlation serve as unit testing of meta-statistical models. However, for more computationally intensive algorithms, such as optimal transport, meta-statistical models could serve as fast approximations. For datasets $\mathcal{D}\in\mathbb{R}^{n\times m}$, we consider four descriptive tasks: the per-column median label $y\in\mathbb{R}^{m}$ consists of the medians of each column. The Pearson correlation coefficient $y\in\mathbb{R}$ is computed between the two columns. The win rate (Bradley-Terry) is the fraction of rows where the value in the first column exceeds that in the second: $y=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(\mathcal{D}_{i,1}>\mathcal{D}_{i,2})$, where $\mathbb{I}(\cdot)$ is the indicator function. Finally, the 1D optimal transport (OT) label $y\in\mathbb{R}$ is the optimal transport cost between the empirical distributions of the two columns.
Meta-Dataset Generation. To construct the meta-dataset, we sample datasets $D$ from predefined probability distributions as described in Section 3. Once a dataset is sampled we simply compute the target label $y$ by applying the target algorithm. We experiment with various numbers of columns $m$. By having $k>1$, we produce $k$ computation in parallel with one forward pass (independently of the batch dimension). We observe no significant difference when varying $k$ and fix $k=2$ in the experiments. The meta-dataset contains $30K$ training meta datapoints per task, with dataset sizes sampled from $n\in[5,300]$. Details about meta-datasets and which distribution families are in- or out-of-meta-distribution are provided in Appendix A.
Meta-Statistical Models. After optimizing hyperparameters and architecture choices (e.g., pooling mechanisms and head-to-dimensionality ratio) on a small validation set of 1K meta datapoints, we compare four meta-statistical model variants: LSTM, Vanilla Transformer (VT), and two ST2 variants with 16 or 32 inducing points. ST2(16) is the fastest model for both training and inference. In Appendix A.4, we show that VT scales quadratically, while LSTM and ST2 scale linearly, with better slopes for ST2. Additionally, ST2(16) achieves a 12x faster training time per batch normalized by parameters compared to VT, meaning an ST2(16) model with 12 times more parameters can be trained in the same time as VT. However, for consistency in reporting, we compare models with approximately the same number of parameters ($\sim 10K$ in this section).
(a) Median prediction
(b) Correlation prediction
Figure 2: Generalization across dataset lengths and meta-distributions. The left panel shows MSE as a function of dataset length for in-meta-distribution datasets, while the right panel displays the same for out-of-meta-distribution datasets. The vertical red line marks the largest dataset seen during training ($n=300$). LSTM is excluded due to its errors being an order of magnitude higher. Additional tasks can be found in Appendix A.3.
In-meta-distribution performance. Table 1 shows the MSE of the four meta-statistical models on a test set sampled from the same meta-distribution as the training data. All models approximate the descriptive tasks well, but the LSTM-based model, lacking permutation invariance, performs worse than attention-based models. Notably, ST2, despite being much faster than VT, narrowly outperforms it. Given its strength and efficiency, ST2(16) is our main model in the rest of the paper, with VT considered as an alternative baseline.
Generalization Performance. We evaluate meta-statistical models’ generalization capabilities on two aspects: (i) Out-of-Meta-Distribution (OoMD): Datasets from unseen distributions. (ii) Length Generalization: Datasets with lengths outside the training range. Figure 2 shows strong length generalization, where models maintain their performance for larger datasets than seen during training, both IMD and OoMD. They are also robust to OoMD datasets despite a small performance degradation. Manual inspection reveals that the degradation mainly comes from cases where the magnitude of the input values exceeds the range seen during training. This is discussed further in Section 6. Additional results and generalization plots are provided in Appendix A.
Median Corr WinRate (BT) OT (1D)
LSTM $2.9e^{-1}$$\pm 0.8$ $5.9e^{-2}$$\pm 1.5$ $4.4e^{-2}$$\pm 0.9$ $8.5e^{-2}$$\pm 2.9$
VT $\mathbf{6.0e^{-2}}$$\pm 1.9$ $\mathbf{9.2e^{-3}}$$\pm 4.6$ $7.1e^{-3}$$\pm 1.5$ $\mathbf{5.5e^{-2}}$$\pm 1.4$
ST2(16) $\mathbf{4.2e^{-2}}$$\pm 1.7$ $\mathbf{7.5e^{-3}}$$\pm 2.8$ $\mathbf{2.9e^{-3}}$$\pm 1.2$ $\mathbf{4.5e^{-2}}$$\pm 1.9$
ST2(32) $\mathbf{4.4e^{-2}}$$\pm 0.9$ $\mathbf{9.1e^{-3}}$$\pm 5.1$ $\mathbf{1.6e^{-2}}$$\pm 0.5$ $\mathbf{3.0e^{-2}}$$\pm 1.5$
Table 1: Performance comparison meta-statistical models across tasks, measured by Mean Squared Error with respect to correct output on the test set. Bold indicates no significant difference with the best model.
5 Experiments on Inferential Tasks
In inferential tasks, the label $y$ represents a property $g$ of the underlying distribution $P_{X}$ from which a dataset $\mathcal{D}$ is sampled: $y=g(P_{X})$. We illustrate the meta-statistical framework with three such tasks: standard deviation estimation, normality testing, and mutual information estimation. Details on meta-dataset creation and models are in Appendix B. For all tasks in this section, the dataset sizes during training are sampled from $n\in[5,150]$, depicted by vertical red lines in the plots.
5.1 Standard Deviation Estimation
The standard deviation ($\sigma=\sqrt{\mathbb{E}[(X-\mathbb{E}[X])^{2}]}$) quantifies the spread of a distribution $P_{X}$. Unlike the mean or variance, estimating $\sigma$ is non-trivial due to the square root’s non-linearity (Gurland and Tripathi, 1971; Gupta, 1952). In fact, no universal unbiased estimator exists across all distributions (Gurland and Tripathi, 1971; Fenstad et al., 1980). We use this task to show meta-statistical learning in action.
Meta-Dataset. To create the meta-dataset, we follow the procedure outlined in Section 3, keeping different distribution families for in- and out-of-meta-distribution. We use 100K meta datapoints for training.
Figure 3: MSE of $\sigma$ estimators as a function of dataset sizes, for dataset sampled out-of-meta-distribution.
[…]
5.2 Normality Testing
The task is now to determine whether a dataset $\mathcal{D}\sim P_{X}$ originates from a normal distribution, formulated as a binary classification task: $y=1$ if $P_{X}$ is normal, $y=0$ otherwise. Normality testing is crucial in hypothesis testing, model selection, and preprocessing (Shapiro and Wilk, 1965; Razali et al., 2011), particularly before applying t-tests, linear regression, or ANOVA with small samples (Altman, 1990; Das and Imon, 2016; Kwak Sang Gyu, 2019). However, standard tests struggle in low-sample settings (Razali et al., 2011). We propose to train meta-statistical models for normality classification, aiming for robust generalization in such regimes. Details on meta-dataset creation and model properties are in Appendix C.
Figure 4: Accuracy of normality classifiers as a function of dataset sizes. The non-normal distributions are sampled out-of-meta distribution for meta-statistical models.
Meta-dataset creation. We construct a balanced meta-dataset of normally and not normally distributed datasets following the process described in Section 3. For non-normality, we choose an alternative distribution from a predefined set detailed in Appendix C. We use 40K meta datapoints for training.
Estimators. We transform traditional normality tests into binary classifiers by thresholding their $p$-values, optimizing the threshold on the training meta-dataset for maximum classification accuracy. We consider four widely used tests: the Shapiro-Wilk test (Shapiro and Wilk, 1965), known to be effective for small samples (Razali et al., 2011); the D’Agostino-Pearson test (D’agostino and Pearson, 1973), which combines skewness and kurtosis; the Kolmogorov-Smirnov test (Massey Jr, 1951), a non-parametric test based on cumulative distribution differences; and the Jarque-Bera test (Jarque and Bera, 1987), which assesses skewness and kurtosis deviations from theoretical expectations.
We then train two meta-statistical models: one based on VT and another on ST2 with 16 inducing points. Both use four layers, a hidden dimensionality of 32, and 12 attention heads. The classification head is a single-layer MLP with 32 neurons, totaling approximately 50K parameters per model.
Accuracy $\uparrow$ AuROC $\uparrow$ Brier $\downarrow$ BT $\uparrow$
KS $0.88$ $0.01$ $0.93$ $0.01$ $0.09$ $0.01$ $0.12$ $0.02$
SW $0.89$ $0.01$ $0.95$ $0.01$ $0.18$ $0.01$ $0.16$ $0.03$
JB $0.88$ $0.01$ $0.93$ $0.01$ $0.16$ $0.01$ $0.13$ $0.02$
AP $0.90$ $0.01$ $0.95$ $0.01$ $0.18$ $0.01$ $0.17$ $0.03$
ST2 $\mathbf{0.92}$ $0.01$ $\mathbf{0.97}$ $0.01$ $\mathbf{0.06}$ $0.01$ $\mathbf{0.25}$ $0.04$
VT $0.91$ $0.01$ $\mathbf{0.97}$ $0.01$ $\mathbf{0.07}$ $0.01$ $0.17$ $0.03$
Table 2: Normality test classifiers with datasets drawn from Gaussian or Uniform distributions of sizes $n\in[10,300]$. AuROC refers to the area under the ROC curve, Brier loss is the calibration error, and BT measures the relative strengths of classifiers in a paired evaluation.
Results. Figure 4 summarizes the accuracy of the proposed meta-statistical models, in settings where negative labels correspond to datasets sampled from distribution families unseen during training. Consistent with prior comparisons of normality tests, Shapiro-Wilk and D’Agostino-Pearson perform best among the baselines (Razali et al., 2011). Across all dataset sizes, meta-statistical models consistently and largely outperform baselines, with particularly strong gains in small-sample settings ($n<100$), making them highly relevant for biomedical applications (Kwak Sang Gyu, 2019). Meta-statistical models achieve near-perfect accuracy ($>0.98$) as $n$ increases demonstrating their consistency. Overall, this task seems relatively easy for meta-statistical models, which generalize smoothly out-of-meta distribution. However, note that the training of meta-statistical models could be harder if the input datasets are standardized during training (see Section 6).
While classification lacks a direct bias-variance formulation, we analyze false positive and false negative rates as well as precision and recall in Appendix C.3, showing more balanced error profiles for meta-statistical estimators. In Table 2, we present key metrics for evaluating classifier performance: the Area Under the Receiver Operating Characteristic Curve (AuROC), the Brier Score, and the Bradley-Terry (BT) scores from a paired evaluation. The AuROC measures a classifier’s ability to discriminate between positive and negative classes across different decision thresholds. A higher AuROC indicates better separability. Unlike accuracy, AuROC provides a threshold-independent measure of performance. The Brier loss (Brier, 1950) quantifies the calibration of a model’s predicted probabilities. Lower values indicate better calibration. The Bradley-Terry (BT) score (Bradley and Terry, 1952; Huang et al., 2004) ranks models based on pairwise comparisons, assessing how often one classifier outperforms another across test instances (Peyrard et al., 2021; Colombo et al., 2023). The accuracy scores are lower than those of Figure 4 because the uniform is among the hardest out-of-meta-distribution to recognize as non-Gaussian. Still, across metrics, the meta-statistical estimators perform strongly. In particular, we find it interesting that they are particularly well-calibrated.
5.3 Mutual Information Estimation
Mutual information (MI) quantifies the dependency between two random variables $X$ and $Y$ and is defined as:
$\mathrm{MI}(X;Y)=\int\int P_{X,Y}(x,y)\log\frac{P_{X,Y}(x,y)}{P_{X}(x)P_{Y}(y)}\,dx\,dy.$
Here, $P_{X}$ and $P_{Y}$ denote the marginal distributions of $X$ and $Y$, respectively.
MI possesses key properties such as invariance to homeomorphisms and adherence to the Data Processing Inequality, making it fundamental in machine learning and related fields (Li et al., 2021; Belghazi et al., 2018; van den Oord et al., 2018; Tishby et al., 2000). However, MI estimation remains challenging, particularly for small sample sizes and non-Gaussian distributions (Song and Ermon, 2020; McAllester and Stratos, 2020; Czyż et al., 2023).
We adopt a meta-statistical approach, training models to predict $y=\mathrm{MI}(X;Y)$ between two dataset columns. Focusing on low-sample, non-Gaussian settings, but we restrict experiments to the one-dimensional case for simplicity. Details on meta-dataset creation, models, and extra results are provided in Appendix D.
Meta-dataset Creation. We construct a meta-dataset inspired by the benchmark methodology in (Czyż et al., 2023), where distributions with ground-truth MI are generated in two steps: (i) by sampling a distribution with known MI, (ii) optionally applying MI-preserving transformations. This process creates complex distributions and datasets with known MI. For generating meta-dataset in this way, we again follow the process described in Section 3 using different base-distribution and MI-preserving transformations between in-meta-distribution and out-of-meta-distribution. We use 50K meta datapoints for training.
Estimators. We compare our approach with the best-performing 1D estimators from (Czyż et al., 2023), including Kraskov-Stögbauer-Grassberger (KSG) (Kraskov et al., 2004), Canonical Correlation Analysis (CCA) (Murphy, 2023), and three neural estimators: MINE (Belghazi et al., 2018), InfoNCE (van den Oord et al., 2018), and NWJE (Nguyen et al., 2007; Nowozin et al., 2016; Poole et al., 2019). We train two meta-statistical models: one based on Vanilla Transformer (VT) and the other on Set Transformer 2 (ST2). Both models consist of five layers, with a hidden dimensionality of 256 and 12 attention heads. The regression head is a single hidden-layer MLP with 128 neurons, resulting in models with approximately 1M parameters.
Estimation Performance. The mean squared error (MSE) results for both in- and out-of-meta-distribution testing are shown in Table 3. Meta-statistical models outperform baseline estimators across all sample sizes, with significant advantages in low-sample scenarios. Baseline models, particularly neural ones, struggle with small sample sizes, while only KSG and CCA begin to match meta-statistical models for sample sizes greater than 100 in the out-of-meta-distribution regime.
Bias and Variance of MI Estimators. We examine the bias and variance of MI estimators by resampling datasets from fixed distributions and measuring the variance and bias of the estimates. In Figure 5, we visualize the bias and variance for a challenging distribution identified by previous works (Czyż et al., 2023) (additive noise). Even at a sample size of $n=100$, meta-statistical models show clear improvements in both bias (estimates centered around 0) and variance. A more detailed analysis of bias and variance is available in Appendix D (Table 7). Compared to baseline estimators, meta-statistical models demonstrate significantly lower bias, close to zero, and lower or comparable variance. These results are promising, suggesting that further scaling could create even more robust meta-statistical MI estimators. Currently, the ST2 model can be trained in less than an hour on a single GPU, with inference orders of magnitude faster than existing neural baselines.
Figure 5: We estimate statistics for MI estimators over 150 resampled datasets of size $n=100$ from a fixed distribution (additive noise (Czyż et al., 2023)). Each dot represents the difference between an estimate and the true mutual information (MI).
IMD OoMD
$n\in$ $[10,100]$ $[100,200]$ $[10,100]$ $[100,200]$
CCA $7.4e^{-2}$$\pm 9.3$ $1.4e^{-2}$$\pm 1.1$ $1.3e^{-1}$$\pm 1.2$ $4.9e^{-2}$$\pm 3.3$
KSG $2.9e^{-2}$$\pm 1.5$ $7.8e^{-3}$$\pm 3.5$ $\mathbf{1.2e^{-2}}$$\pm 0.4$ $\mathbf{7.2e^{-3}}$$\pm 2.3$
MINE $2.5e^{0}$$\pm 2.4$ $2.8e^{-2}$$\pm 1.2$ $5.4e^{0}$$\pm 7.7$ $1.6e^{-1}$$\pm 2.1$
NWJE - $7.3e^{-2}$$\pm 4.7$ $6.6e^{0}$$\pm 7.5$ $6.3e^{-2}$$\pm 5.4$
InfoNCE $1.5e^{1}$$\pm 2.2$ $1.9e^{-2}$$\pm 0.7$ $2.3e^{1}$$\pm 3.4$ $3.4e^{-1}$$\pm 4.5$
VT $\mathbf{4.6e^{-3}}$$\pm 2.4$ $\mathbf{2.5e^{-3}}$$\pm 1.3$ $\mathbf{1.5e^{-2}}$$\pm 0.8$ $\mathbf{7.7e^{-3}}$$\pm 3.2$
ST2(16) $\mathbf{6.2e^{-3}}$$\pm 3.0$ $\mathbf{2.4e^{-3}}$$\pm 1.1$ $\mathbf{1.3e^{-2}}$$\pm 0.7$ $\mathbf{8.5e^{-3}}$$\pm 3.1$
Table 3: MSE loss of mutual information estimators both in- and out-of-meta-distribution. Bold indicates no significant difference with the best estimator.
6 Discussion
With the meta-statistical framework, statistical inference becomes synonymous with inference in machine learning, and what is hard for statistical inference reveals itself as hard to learn for our models. Our experiments reveal interesting difficulties in statistical inference. Predicting normality was the easiest task for meta-statistical models, requiring only 50K parameters for strong generalization. Estimating mutual information, as expected, demanded significantly larger models ( 1M parameters). Surprisingly, predicting the standard deviation was particularly difficult: while small models ( <10K parameters) could easily approximate sample standard deviation (descriptive), nearly 1M parameters were needed to predict the standard deviation (inferential) better than np.std. Training a model to predict only the corrective term also required nearly 1M parameters and yielded an estimator equivalent to directly estimating the true standard deviation, suggesting that finite-sample errors is the main driver of difficulty in this problem. This raises intriguing questions about what makes meta-statistical models work and fail: do these models implicitly perform Bayesian inference with input-dependent priors?
Limitations and Future Work. While meta-statistical learning brings the advantages of machine learning to statistical inference, it also imports its challenges. A key question is the choice of meta-distribution during training—what constitutes a good meta-distribution to sample from? Additionally, the evaluation of estimators becomes more difficult; a model trained on a narrow meta-distribution might generalize poorly outside its training regime.
Interpretability is another challenge. The precise algorithm computed at inference to perform the statistical inference becomes unknown and difficult to interpret (Teney et al., 2022). Also, like LLMs, meta-statistical models could exhibit unexpected failure cases and lack strict guarantees of validity. However, they also offer a promising testbed for mechanistic interpretability research: they process structured numerical inputs without tokenization, operate in a single forward pass, and construct mathematical representations rather than linguistically ambiguous ones.
Failure cases also merit further study. Models struggled when input scales exceeded training ranges and we found one case of poor generalization to one unseen distribution family (log-normal) in the standard deviation estimation task (documented in Appendix B.4). In the normality test setting, we believe that standardizing the datasets would make training harder but encourage better generalization OoMD by preventing the meta-statistical estimators from picking up on spurious associations between the meta-distribution and the labels. Overall, like LLMs, these models would benefit from larger and more diverse training data. Future directions include learned row embeddings to accommodate varying input row dimensions and magnitudes, as well as scaling laws to guide the training of larger models with optimized data mixes. One limitation of this work is the focus on one-dimensional datasets to explore inference tasks in a controlled setting, but real-world inference involves high-dimensional data, where traditional estimators often struggle. Scaling to higher dimensions is a key next step, and we anticipate meta-statistical models to generalize well. In general, this work focuses on demonstrating promises of the meta-statistical perspective encouraging further efforts in crafting and evaluating learned statistical estimators with methods inspired by natural language processing.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. Meta-statistical learning aims to enhance inference in low-sample settings, benefiting applied fields of Science like medicine and economics by improving estimator reliability. Learned estimators may inherit biases from the data they are trained on, potentially leading to misleading conclusions if not carefully validated. Further, as with any data-driven methodology, interpretability remains a challenge; understanding why a model makes a particular statistical inference is crucial for scientific rigor.
Appendix A Details about the Descriptive tasks experiments
A.1 Details of Meta-Dataset Creation
To ensure reproducibility of the experiments, we describe the synthetic data generation process in detail. The datasets were generated using a custom-built class, DescMetaDatasetGenerator, which allows for the creation of datasets with various distributions and customizable descriptive target variables. The key components and configurations are outlined below.
In-Meta-Distribution. The set of distributions used to generate datasets during training is parameterized as follows:
  • normal:: It has two parameters: the mean and the variance. Mean values are sampled from [-3, 3], and variances are sampled from [0.1, 1.5].
  • uniform:: It has two parameters: the lower bound and the upper bound. The lower bounds are sampled from [-3.5, -0.5] and the upper bounds from [0.5, 3.5].
  • beta:: It has two parameters: $a$ and $b$. Parameters $a$ and $b$ are sampled from [1, 3] and [2, 5], respectively.
  • exponential:: It has one parameter: scale sampled from [1, 2].
Out-of-Meta-Distribution. The set of distribution used to test models for unseen distribution families is parametrized as follows:
  • gamma:: It has two parameters: shape and scale. Shape parameters are sampled from [1, 5], and scale parameters from [1, 2].
  • log-normal:: It has two parameters: mean and variance. Means are sampled from [0, 1], and standard deviations from [0.5, 0.75].
Dataset Characteristics. Once a distribution $P_{X}$ has been sampled, we use it to sample one dataset. In general, we could sample several dataset per distributions but we prefer to sample only one to maximize the diversity of distributions seen during training. Each dataset is defined by the following parameters:
[…]
  • Number of rows (n_row_range): The number of samples (rows) in the dataset, sampled uniformly from the range [5, 300]. During testing, we explore longer lengths to test the generalization of meta-statistical models.
  • To generate the target values $y$, each dataset is passed through the target descriptive functions: per-column-mean, per-column-median, correlation, win rate, optimal transport (1D).
This results in a meta-datapoint. We them sample many meta-datapoints to build a meta-datasets with the following split sizes: 30K training, 300 validation, 3K for testing in-meta-distribution and 3K for testing out-of-meta-distribution.
A.2 Examples of Training Curves
For meta-statistical models of approximately the same size ($\approx$ 10K parameters), we compare their convergence during training on the task of predicting the correlation between variable A and variable B, the two columns of the dataset. We consider the same meta-statistical models and the same meta-dataset generation parameters as considered in the results of the main paper. We report the results in Figure 6.
Figure 6: Training curves: Comparison of training convergence of meta-statistical models on the correlation task.
A.3 More Generalization Plots
In Figure 7, we report the same generalization plots as the main paper for other descriptive tasks. Similar conclusion holds: models generalize very well with lengths and tend to suffer from an offset of performance out-of-meta-distribution.
(a) Win-rate prediction
(b) Median prediction
(c) Correlation prediction
(d) Optimal Transport prediction
Figure 7: Generalization Across Dataset Lengths and Meta-Distributions. For each subplot, the left panel illustrates the performance of meta-statistical models on test datasets that vary in input length, including lengths not observed during training, while remaining within the training meta-distribution. For each subplot, the right panel presents the same comparison but for test datasets sampled from entirely new meta-distributions, with distributions unseen during training. Note that LSTM is excluded because its errors are an order of magnitude higher.
A.4 Details about Efficiency
In Figure 8, we present the inference time of meta-statistical models as a function of the input dataset size $n$. As expected, the VT scales quadratically, whereas LSTM and ST2 variants scale linearly with slopes in favor of ST2. We also compare the efficiency per parameter. For this we compute both the training and inference time of each model per batch averaged over 1K batches, and normalized by the number of parameters in the model. The results are reported in Table 4. Given the strong performance of ST2 and the clear computational advantage we see it as strong meta-statistical architecture.
Figure 8: Inference time comparison of meta-statistical models per batch as a function of input dataset length. Models have similar parameter counts $\approx 10K$).
Training Time Inference Time
VT $2.2e^{-5}$$\pm 1.1$ $2.7e^{-3}$$\pm 1.2$
LSTM $6.5e^{-6}$$\pm 1.0$ $8.5e^{-4}$$\pm 3.9$
ST2(32) $5.9e^{-6}$$\pm 6.9$ $1.5e^{-3}$$\pm 2.4$
ST2(16) $1.7e^{-6}$$\pm 0.2$ $2.2e^{-4}$$\pm 0.9$
Table 4: Comparison of training and inference times for various models, normalized per batch per number of parameters.
Appendix B Details about Standard Deviation Experiments
B.1 Details of Meta-Dataset Creation
We construct a meta-dataset by generating datasets labeled with the ground truth standard deviation, using a set of distributions for which the standard deviation is well-defined. To create each meta-datapoint, we first sample the base distribution uniformly at random from a set of pre-defined distribution families (see below). Then, we sample the parameters of the distribution, resulting in a distribution $P_{X}$. A dataset size $n$ is then drawn uniformly at random from the range $[10,150]$, and the dataset $D$ is sampled with $n$ rows. We generate 50K meta-datapoints for training and 3K for validation.
In-Meta-Distribution. These are the distributions seen during training. The base distributions are the following, with the priors on their parameters:
  • normal: the mean is sampled from $\mathcal{U}(-1,1)$, and the variance is sampled from $\mathcal{U}(0.5,2.0)$.
  • uniform: the lower bound is sampled from $\mathcal{U}(0,0.5)$, and the upper bound is sampled from $\mathcal{U}(0.5,1.5)$.
  • exponential: the scale parameter is sampled from $\mathcal{U}(1,2)$.
  • gamma: the shape parameter is sampled from $\mathcal{U}(1,5)$, and the scale parameter is sampled from $\mathcal{U}(1,2)$.
Out-of-Meta-Distribution. These are the distributions seen during training. The base distributions are the following, with the priors on their parameters:
  • beta: the $\alpha$ parameter is sampled from $\mathcal{U}(1,5)$, and the $\beta$ parameter is sampled from $\mathcal{U}(1,5)$.
  • lognormal: the mean of the underlying normal distribution is sampled from $\mathcal{U}(0,1)$, and the standard deviation from $\mathcal{U}(0.1,1)$.
  • weibull: the shape parameter is sampled from $\mathcal{U}(1,5)$, and the scale parameter is sampled from $\mathcal{U}(1,2)$.
B.2 Details about Meta-Statistical Models
For these experiments, we train two meta-statistical models based on Set Transformer 2 (ST2). Variants with different numbers of inducing points were tested, such as ST2(16), which uses 16 inducing points (num_inds = 16).
[…]
Training Configuration. The models were trained on a regression task using the binary MSE loss. We employed a batch size of 64 and optimized the model parameters using the Adam optimizer with a learning rate of $1\times 10^{-5}$. The training process spanned 10 epochs.
B.3 Bias and Variance
In Figure 9, we report an experiment targeted at measuring bias and variance by resampling 150 datasets from a fixed exponential distribution. Let the true standard deviation be denoted as $\sigma$ and the estimates for the $i$-th dataset be $\hat{\sigma}_{i}$, for $i=1,\dots,150$.
The bias of the estimator is computed as:
$\text{Bias}=\frac{1}{n}\sum_{i=1}^{n}\hat{\sigma}_{i}-\sigma,$ (1)
where $n=50$ is the number of resampled datasets.
The variance of the estimator is computed as:
$\text{Variance}=\frac{1}{n}\sum_{i=1}^{n}\left(\hat{\sigma}_{i}-\frac{1}{n}\sum_{j=1}^{n}\hat{\sigma}_{j}\right)^{2}.$ (2)
It shows that the meta-statistical models can improve over the sample standard deviation np.std both in terms of bias and variance. Especially, the learned correction is capable of reducing the bias of np.std indicating that it has not just learned a constant offset. We also report bias, variance, and MSE across dataset sizes and over many meta-datapoints sampled from various distributions (both in- and out-of-meta-distribution) in Table 5.
Figure 9: Estimate statistics for standard deviation estimators over 150 resampled datasets of size $n\in[10,50]$ for the exponential distributions. Each dot represents the difference between an estimate and the true standard deviation. An unbiased estimator should be centered around zero.
Bias Variance MSE
$n\in[10,50]$ $n\in[50,100]$ $n\in[100,150]$ $n\in[10,50]$ $n\in[50,100]$ $n\in[100,150]$ $n\in[10,50]$ $n\in[50,100]$ $n\in[100,150]$
np.std $-3.4e^{-2}$$\pm 2.3$ $-2.5e^{-2}$$\pm 1.6$ $-3.3e^{-3}$$\pm 14.3$ $1.1e^{-1}$$\pm 0.3$ $3.5e^{-2}$$\pm 0.9$ $1.7e^{-2}$$\pm 0.4$ $1.2e^{-1}$$\pm 0.2$ $4.3e^{-2}$$\pm 0.6$ $2.2e^{-2}$$\pm 0.3$
ST2std $-8.1e^{-3}$$\pm 27.7$ $1.9e^{-2}$$\pm 2.1$ $1.8e^{-2}$$\pm 1.4$ $1.6e^{-2}$$\pm 0.3$ $1.1e^{-2}$$\pm 0.2$ $6.4e^{-3}$$\pm 1.2$ $3.7e^{-2}$$\pm 0.4$ $2.2e^{-2}$$\pm 0.3$ $1.1e^{-2}$$\pm 0.1$
np.std + ST2fsd $-5.4e^{-3}$$\pm 28.6$ $-1.5e^{-2}$$\pm 2.2$ $-1.1e^{-2}$$\pm 1.4$ $3.6e^{-2}$$\pm 1.0$ $1.5e^{-2}$$\pm 0.4$ $8.4e^{-3}$$\pm 1.8$ $5.7e^{-2}$$\pm 0.8$ $2.5e^{-2}$$\pm 0.3$ $1.4e^{-2}$$\pm 0.2$
np.std $-5.6e^{-2}$$\pm 2.9$ $-3.7e^{-2}$$\pm 2.1$ $-1.4e^{-2}$$\pm 1.6$ $1.7e^{-1}$$\pm 1.1$ $7.9e^{-2}$$\pm 4.9$ $1.9e^{-2}$$\pm 1.0$ $2.0e^{-1}$$\pm 0.8$ $9.1e^{-2}$$\pm 4.3$ $2.6e^{-2}$$\pm 0.7$
ST2std $7.1e^{-2}$$\pm 9.9$ $-2.1e^{-2}$$\pm 6.5$ $1.2e^{-2}$$\pm 5.2$ $1.3e^{-2}$$\pm 0.3$ $8.0e^{-3}$$\pm 1.9$ $5.0e^{-3}$$\pm 1.6$ $2.7e^{-1}$$\pm 0.3$ $1.1e^{-1}$$\pm 0.2$ $7.9e^{-2}$$\pm 1.5$
np.std + ST2fsd $4.9e^{-2}$$\pm 6.0$ $-4.8e^{-2}$$\pm 4.3$ $-3.2e^{-2}$$\pm 3.4$ $9.1e^{-2}$$\pm 8.3$ $4.8e^{-2}$$\pm 4.0$ $1.0e^{-2}$$\pm 0.6$ $1.9e^{-1}$$\pm 0.5$ $9.8e^{-2}$$\pm 4.2$ $4.0e^{-2}$$\pm 1.1$
Table 5: Bias and Variance of Standard Deviation Estimators The top part shows the in-meta-distribution results, while the bottom part reports the out-of-meta-distribution results.
B.4 Interesting Failure Case
While meta-statistical models generally demonstrate strong robustness when presented with datasets sampled from distribution unseed during training (OoMD), we found a interesting failure case. For the log-normal family that was unseen during training, meta-statistical estimators of standard deviations failed to provide improvements over np.std and even performed worse. This is illustrated in Figure 10 across dataset sizes. Log-normal is a skewed distribution which is particularly challenging estimators of standard deviation which can explain why meta-statistical model poorly generalize in this case. These models can benefit from a larger and more diverse training meta-dataset to exhibit even more robust generalization. Note, however, that the models do not fail for other unseen distribution families as it can be seen in Figure 3 of the main paper.
Figure 10: MSE of $\sigma$ estimators as a function of dataset sizes, for the log-normal distribution.
Appendix C Details about Normality Tests Experiments
C.1 Details of Meta-Dataset Creation
We construct a meta-dataset by generating datasets labeled with the ground truth binary indicator of normality, using a diverse set of alternative distributions. In previous studies, the uniform distribution was used the contrast distribution Razali et al. (2011). To create each meta-datapoint, we first determine whether the dataset is sampled from a normal distribution by flipping a fair coin. If the outcome is normal, we sample the mean and variance of the Gaussian distribution; otherwise, we sample uniformly at random one distribution from the set of non-Gaussian distributions and then draw its parameters.. A dataset size $n$ is then drawn uniformly at random from the range $[5,150]$, and the dataset $D$ is sampled with $n$ rows. We generate 40K meta-datapoints for training, 1K for validation, and 1K meta-datapoints for testing.
The normal distribution parameters are sampled according to: the mean is sampled from $\mathcal{U}(-3,3)$, and the standard deviation from $\mathcal{U}(0.5,3)$.
In-Meta-Distribution. These are the distributions seen during training as non-normal distributions. Not however that because the distribution sampled parameters and the datasets are then sampled from the distributions, in-meta-distribution evaluation still produces different meta-datapoints.
  • gamma: the shape parameter is sampled from $\mathcal{U}(1,5)$, and the scale parameter from $\mathcal{U}(0.5,2)$.
[…]
  • cauchy: the location parameter is sampled from $\mathcal{U}(-1,1)$, and the scale parameter from $\mathcal{U}(0.5,2)$.
  • laplace: the location parameter is sampled from $\mathcal{U}(-1,1)$, and the scale parameter from $\mathcal{U}(0.5,2)$.
  • weibull: the shape parameter is sampled from $\mathcal{U}(0.5,5)$.
  • vonmises: the mean direction $\mu$ is sampled from $\mathcal{U}(-\pi,\pi)$, and the concentration parameter $\kappa$ from $\mathcal{U}(0.5,5)$.
  • arcsine: the lower bound is sampled from $\mathcal{U}(-3,0)$, and the upper bound from $\mathcal{U}(0,3)$. Data is generated by transforming uniform samples with a sine function scaled to the specified bounds.
  • bimodal: two Gaussian components are used, where the means of the components are sampled from $\mathcal{U}(-3,-1)$ and $\mathcal{U}(1,3)$, respectively. The standard deviations are sampled from $\mathcal{U}(0.5,1)$, and the mixture ratio is sampled from $\mathcal{U}(0.3,0.7)$.
Out-of-Meta-Distribution. These are the distributions unseen during training as non-normal distributions.
  • uniform: the lower bound is sampled from $\mathcal{U}(-3,0)$, and the upper bound from $\mathcal{U}(0,3)$.
  • exponential: the scale parameter is sampled from $\mathcal{U}(0.5,2)$.
  • beta: the shape parameters $a$ and $b$ are sampled independently from $\mathcal{U}(0.5,5)$.
  • log-normal: the mean is sampled from $\mathcal{U}(-1,1)$, and the standard deviation from $\mathcal{U}(0.5,1.5)$.
C.2 Details of Meta-Statistical Models
The experiments compared the performance of multiple model variants, including:
  • Vanilla Transformer (VT): This model utilized the vanilla_transformer architecture.
  • Set Transformer 2 (ST2): a set Transformer with 16 inducing points (num_inds = 16).
Both models have four layers, hidden dimensionality of 32, and 12 attention heads. To build a full meta-statistical model from the meta-statistical encoders, we add a prediction head made of a MLP with one layer of 32 neurons before predicting the probability of being normally distributed. The meta-model based on VT has a total of $51K$ parameters, and the meta-model based on ST2 has a total of $54K$. ST2 has more parameters to learn the projected attention but ends being much faster to train and use at inference because of the constant cost of attention compared to quadratic in the length of the input for VT.
Training Configuration. The models were trained on a binary classification task using the binary cross-entropy loss. We employed a batch size of 24 and optimized the model parameters using the Adam optimizer with a learning rate of 0.0005. The training process spanned 7 epochs. These hyper-parameters were selected to balance computational efficiency and convergence stability. In the main paper Figure 4, the meta-statistical models are evaluated OoMD. For completeness, we report their results on unseen test set but in-meta-distribution in Table 11
Figure 11: Accuracy of meta-statistical models compared to standard normality tests converted into classifiers using optimized $p$-value thresholds. The non-normal distributions are sampled in-meta-distribution for meta-statistical models.
C.3 Details about Precision and Recall
While classification lacks a direct bias-variance formulation, we analyze false positive and false negative rates as well as precision and recall in Table 6. Interestingly, most baselines, when optimizing their p-value threshold for maximum accuracy end up maximizing recall at the expense of precision with high false positive rates. On the contrary, the meta-statistical models perform well in both precision and recall with balanced error profiles, i.e., similar amount of false negatives and false positives.
TP FP TN FN Prec. Recall F1-Score
Shapiro-Wilk 439 138 383 40 0.7608 0.9165 0.8314
Kolmogorov-Smirnov 479 364 157 0 0.5682 1.0000 0.7247
D’Agostino and Pearson 423 147 374 56 0.7421 0.8831 0.8065
Jarque Bera 449 296 225 30 0.6027 0.9374 0.7337
VT 471 38 441 50 0.9253 0.9040 0.9146
ST2(16) 474 50 429 47 0.9046 0.9098 0.9072
Table 6: Reporting True Positive (TP), False Positive (FP), True Negative (TN), False Negative (FN), the associated Precision, Recall, and F1-Score of normality classifers over 1,000 datasets of size $n=50$.
Appendix D Details about Mutual Information Experiments
D.1 Details of Meta-Dataset Creation
We construct a meta-dataset inspired by the benchmark methodology in (Czyż et al., 2023), where distributions with ground-truth MI are generated in two steps: (i) by sampling a distribution with known MI, (ii) potentially applying MI-preserving transformation. This process creates complex distributions and datasets with known MI. For generating meta-dataset in this way, we again follow the process described in Section 3 using different base-distribution and MI-preserving transformation between in-meta-distribution and out-of-meta-distribution.
In-Meta-Distribution. The in-meta-distributions focus on bivariate relationships and transformations:
  • binormal-[base, wigglify, halfcube, asinh, normal_cdfise]: Standard bivariate normal distribution with correlation sampled from $\mathcal{U}(-1,1)$, following by any of the MI-preserving transformation (or none).
  • bimodal_gaussians-base: Bivariate Gaussian mixture model with correlation sampled from $\mathcal{U}(-1,1)$.
  • bistudent-[base, asinh]: Bivariate Student’s t-distribution with degrees of freedom sampled from $\mathcal{U}(1,10)$, potentially followed by asinh transform.
Out-of-Meta-Distribution. The out-of-meta-distributions extend the variety of data by incorporating additional transformations and configurations:
  • additive_noise-[base, wigglify, halfcube, asinh, normal_cdfise]: A model where $X$ is sample from $\mathcal{U}(-1,1)$ and $Y=X+\epsilon$, a Gaussian noise sampled from $\mathcal{U}(0.01,2)$) followed by any of the MI-preserving transformation.
  • bistudent-[wigglify, halfcube, normal_cdfise]: Bivariate Student’s t-distribution (degrees of freedom sampled from $\mathcal{U}(1,10)$) followed by any of the MI-preserving transformation except asinh.
For the computation of the true MI from these distributions and their transformations, we refer to (Czyż et al., 2023); we sample using the tool they provided: https://github.com/cbg-ethz/bmi.
D.2 Details of the Training of Meta-Statistical Models
The meta-statistical models were trained using the code framework outlined in the previous section. This section provides detailed information about the training process.
Default Configuration. The default configuration specifies the following key components:
  • Dimensionality: In these experiments, we focus on 1D dimensional variables.
  • Dataset Parameters: Dataset sizes are sampled uniformly from the range [10, 150].
  • Meta-Dataset Properties: The meta-dataset contains 50,000 training, 500 validation, and 1,000 testing meta-datapoints.
  • Training Parameters: A regression task was specified with a batch size of 64, learning rate of 0.0001, and 20 epochs of training.
  • Model Architecture: The base model used a 5-layer encoder (n_enc_layers), with hidden dimensions of 256 and 128 for the phi and theta components, respectively. Multi-head attention mechanisms used 12 heads.
Model Variants. The experiments compared the performance of multiple model variants, including:
  • Vanilla Transformer (VT): This model utilized the vanilla_transformer architecture.
  • Set Transformer 2 (ST2): a set Transformer with 16 inducing points (num_inds = 16).
To build a full meta-statistical model from the meta-statistical encoders, we add a prediction head made of a MLP with one layer of 128 neurons before predicting the 1 number target output. The meta-model based on VT has a total of $1,008,385$ parameters, and the meta-model based on ST2 has a total of $1,280,065$. ST2 has more parameters to learn the projected attention but ends being much faster to train and use at inference because of the constant cost of attention compared to quadratic in the length of the input for VT.
D.3 Details about the Bias and Variance of MI Estimators
To estimate the bias and variance of a mutual information (MI) estimator, we follow a systematic procedure. First, we sample a base distribution and an MI-preserving transformation from the out-of-meta-distribution set. From this fixed setup, we resample $n=50$ datasets, leading to 50 MI estimates from the given estimator. Let the true mutual information be denoted as $I_{\text{true}}$ and the MI estimates for the $i$-th dataset be $\hat{I}_{i}$, for $i=1,\dots,50$.
The bias of the estimator is computed as:
$\text{Bias}=\frac{1}{n}\sum_{i=1}^{n}\hat{I}_{i}-I_{\text{true}},$ (3)
where $n=50$ is the number of resampled datasets.
The variance of the estimator is computed as:
$\text{Variance}=\frac{1}{n}\sum_{i=1}^{n}\left(\hat{I}_{i}-\frac{1}{n}\sum_{j=1}^{n}\hat{I}_{j}\right)^{2}.$ (4)
To capture a broader picture of estimator behavior, we repeat this process for 100 random choices of base distributions and transformations. For each random choice, we report the bias and variance of the estimator as calculated above. Finally, we summarize these results across different sample sizes in Table 7.
Bias Variance
$n\in$ $[10,50]$ $[50,100]$ $[100,150]$ $[10,50]$ $[50,100]$ $[100,150]$
CCA $-5.0e^{-2}$$\pm 3.7$ $-1.5e^{-2}$$\pm 1.2$ $-3.7e^{-2}$$\pm 1.8$ $8.6e^{-3}$$\pm 3.3$ $3.5e^{-3}$$\pm 1.2$ $1.3e^{-3}$$\pm 0.5$
KSG $-1.0e^{-1}$$\pm 0.6$ $-2.6e^{-2}$$\pm 2.9$ $-2.9e^{-2}$$\pm 1.2$ $5.6e^{-3}$$\pm 1.5$ $5.6e^{-3}$$\pm 1.7$ $1.8e^{-3}$$\pm 0.7$
VT $4.4e^{-3}$$\pm 10.8$ $4.0e^{-3}$$\pm 9.7$ $-7.8e^{-4}$$\pm 57.2$ $9.8e^{-3}$$\pm 7.1$ $3.1e^{-3}$$\pm 1.2$ $1.1e^{-3}$$\pm 0.5$
ST2(16) $1.3e^{-2}$$\pm 1.0$ $9.8e^{-3}$$\pm 9.8$ $7.5e^{-4}$$\pm 60.7$ $7.7e^{-3}$$\pm 4.5$ $3.6e^{-3}$$\pm 1.4$ $9.8e^{-4}$$\pm 3.9$
Table 7: Bias and Variance of Mutual Information Estimators
Bias Variance MSE
$n\in[10,50]$ $n\in[50,100]$ $n\in[100,150]$ $n\in[10,50]$ $n\in[50,100]$ $n\in[100,150]$ $n\in[10,50]$ $n\in[50,100]$ $n\in[100,150]$
np.std $-1.0e^{-2}$$\pm 1.1$ $1.3e^{-2}$$\pm 0.2$ $-2.2e^{-2}$$\pm 0.7$ $1.9e^{-2}$$\pm 0.4$ $8.1e^{-3}$$\pm 1.6$ $3.8e^{-3}$$\pm 0.6$ $2.0e^{-2}$$\pm 0.2$ $8.2e^{-3}$$\pm 0.6$ $3.9e^{-3}$$\pm 0.3$
ST2std $-1.2e^{-1}$$\pm 0.8$ $2.2e^{-3}$$\pm 9.7$ $-2.4e^{-2}$$\pm 0.9$ $8.6e^{-3}$$\pm 1.3$ $5.1e^{-3}$$\pm 0.6$ $3.4e^{-3}$$\pm 0.5$ $1.5e^{-2}$$\pm 0.1$ $7.5e^{-3}$$\pm 0.5$ $3.9e^{-3}$$\pm 0.3$
ST2np.std +ST2fsd $2.0e^{-2}$$\pm 0.8$ $4.5e^{-2}$$\pm 1.6$ $9.2e^{-3}$$\pm 8.5$ $8.9e^{-3}$$\pm 1.6$ $5.7e^{-3}$$\pm 1.0$ $3.7e^{-3}$$\pm 0.6$ $1.4e^{-2}$$\pm 0.1$ $6.9e^{-3}$$\pm 0.5$ $3.9e^{-3}$$\pm 0.3$
np.std + ST2fsd $1.4e^{-3}$$\pm 12.0$ $3.4e^{-2}$$\pm 0.7$ $4.7e^{-3}$$\pm 4.6$ $1.0e^{-2}$$\pm 0.1$ $5.8e^{-3}$$\pm 0.7$ $3.6e^{-3}$$\pm 0.5$ $1.4e^{-2}$$\pm 0.1$ $6.9e^{-3}$$\pm 0.5$ $3.6e^{-3}$$\pm 0.3$
Table 8: Bias and Variance of Standard Deviation Estimators